OpenStudy (anonymous):

If x= y-3 and y=z^2, what is x in terms of z? Answer: a. z^-6 b. z^6 c. z^-2/3 d. z^5 e. z^3/2 i keep getting x=z^2-3

OpenStudy (anonymous):

x=y-3 x=z^2 -3

OpenStudy (anonymous):

I guess x=y^(-3)

OpenStudy (anonymous):

Can u PLZ check it

OpenStudy (anonymous):

check it how? i copy and pasted the problem exactly how it was displayed , even if x=y^-3 the answer would be x=-z and that isnt an option ethier

OpenStudy (anonymous):

If x=y^-3 then x=(z^2)^-3 =z^-6

OpenStudy (anonymous):

oh, i see how that works, thanks , it must be x=y^3 , it had another problem like that earlier where it didn't show like a exponent but it was.